OpenStudy (anonymous):

How long it will take for $300 to double when invested at 6% annual interest compounded twice a year? 117 Which of the following is a solution to the equation log 6 x + log 6 (x + 1) = 1 ? 2

OpenStudy (anonymous):

A painting is purchased for $350. If the value of the painting doubles every 5 years, then its value is given by the function V(t) = 350 • 2t/5, where t is the number of years since it was purchased and V(t) is its value (in dollars) at that time. What is the value of the painting ten years after its purchase? 1,400 What is the solution to the equation 21x + 3 = 14 x ≈ –2.133

OpenStudy (anonymous):

What is the solution to the equation 7^ 3x ≈ 9 x = 0.376

OpenStudy (anonymous):

are those all right so far ?

ganeshie8 (ganeshie8):

hold on a bit

ganeshie8 (ganeshie8):

first q, you mean Answer is 11.7 years ?

OpenStudy (anonymous):

yesss .

ganeshie8 (ganeshie8):

first q = \(\color{red}{\checkmark}\)

OpenStudy (anonymous):

(; . okkk good

ganeshie8 (ganeshie8):

second q = \(\color{red}{\checkmark}\)

OpenStudy (anonymous):

(:

ganeshie8 (ganeshie8):

for 3rd q, 1400 = \(\color{red}{\checkmark}\)

OpenStudy (anonymous):

(:

ganeshie8 (ganeshie8):

for 4th q, x ≈ –2.133 = \(\color{red}{\checkmark}\)

OpenStudy (anonymous):

okkk . (: are you ready for the nesxt few ?

ganeshie8 (ganeshie8):

5th q also \(\color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

yea you seem to be always right lol

OpenStudy (anonymous):

hahaha (:. Which of the following is equivalent to log 40 8 rounded to three decimal places? 0.564 What is the solution of log 2x + 3125 = 3 ? x = 4 Which of the following is equivalent to log 2x 3y ? log(3y)/log(2x) Which of the following is the simplified version of 3 log x + log y + log z ? logx^3y/z

ganeshie8 (ganeshie8):

0.564 is correct

ganeshie8 (ganeshie8):

but next question is wrong

OpenStudy (anonymous):

):. awee ok . lets see what i did wrong

ganeshie8 (ganeshie8):

What is the solution of \(\log_{ 2x + 3}125 = 3\) ?

ganeshie8 (ganeshie8):

is it really 2x+3, or 2x-3 ?

OpenStudy (anonymous):

2x+3,

ganeshie8 (ganeshie8):

cuz i remember, we did the 2x-3 in subscript before

ganeshie8 (ganeshie8):

just check once

OpenStudy (anonymous):

yup but thay changed the subscript

ganeshie8 (ganeshie8):

ohk... lol then answer changes too

ganeshie8 (ganeshie8):

lets solve it real quick

OpenStudy (anonymous):

okkk (:

ganeshie8 (ganeshie8):

\(\log_{ 2x + 3}125 = 3 \) \(125 = (2x+3)^3\) \(5^3 = (2x+3)^3\) \(5 = 2x + 3\) solve x

OpenStudy (anonymous):

2=2x x=1 okk. i see what i did wrong i mad it 8=2x on accident

ganeshie8 (ganeshie8):

ok good, next

ganeshie8 (ganeshie8):

next q is \(\color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

last q is wrong. they changed the q again it seems

ganeshie8 (ganeshie8):

Which of the following is the simplified version of \(3 \log x + \log y + \log z\) ?

ganeshie8 (ganeshie8):

you sure, all + signs u have there ?

OpenStudy (anonymous):

oh sorry no i had put logx^3/yz. but we can be sure and check it .

ganeshie8 (ganeshie8):

\(3 \log x + \log y + \log z\) all + signs ? can u look at q and confirm

OpenStudy (anonymous):

yess thats right.

ganeshie8 (ganeshie8):

then u simply multiply

ganeshie8 (ganeshie8):

\(3 \log x + \log y + \log z\) \(\log x^3 + \log y + \log z\) \(\log x^3 y z\)

OpenStudy (anonymous):

so ii was right ?

ganeshie8 (ganeshie8):

cuz \(\log a + \log b = \log ab\)

ganeshie8 (ganeshie8):

Which of the following is the simplified version of 3 log x + log y + log z ? logx^3y/z

ganeshie8 (ganeshie8):

you have placed z in the bottom, so you're not right

OpenStudy (anonymous):

ohhhhhhhhhhhhh! okkkk . sorry lol i got it now.

OpenStudy (anonymous):

ok last few. What is the solution to the equation 9(x + 1) = 27 ? x = 0.5 What is the solution to the equation 1√8=4^(m + 3) ? -11/4 The value of a dirt bike decreases by 15% each year. If you purchased this dirt bike today for $500, to the nearest dollar how much would the bike be worth 5 years later? 222

ganeshie8 (ganeshie8):

its okay :)

OpenStudy (anonymous):

(:

ganeshie8 (ganeshie8):

first q, \(\color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

second q, wrong

OpenStudy (anonymous):

ok im going to redo it.

ganeshie8 (ganeshie8):

let me check 3rd q, before coming back to second q

ganeshie8 (ganeshie8):

ok you try, meantime il check 3rd q ok

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

for last q, 222 = \(\color{red}{\checkmark}\)

OpenStudy (anonymous):

ok the last one is Which of the following is equivalent to log 4 32 ? 2.5

ganeshie8 (ganeshie8):

thats right, wat about the wrong q

OpenStudy (anonymous):

ok i redid the second one and got -15/4. im not sure its right though.

ganeshie8 (ganeshie8):

-15/4 = \(\color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

Excellent work !

OpenStudy (anonymous):

yayyy thank you veryyy muchhh (:

ganeshie8 (ganeshie8):

ou got more

OpenStudy (anonymous):

yes i do . im going to run an arron then tagg you in any more . btw i got a 100% (:

ganeshie8 (ganeshie8):

congrats !! you deserve 100 %